Find the slope-intercept equation of the line passing through (1,2) with a slope of m=3
Found 2 solutions by stanbon, Aldorozos:
Answer by stanbon(75887)

Find the slope-intercept equation of the line passing through (1,2)
with a slope of m=3
----------------
Form:: y = mx + b
Solve for "b" when m = 3, x = 1, y = 2
------------------------
2 = 3*1 + b
b = -1
------
Equation:
y = 3x -1
----------------
Cheers,
Stan H.
------------

Answer by Aldorozos(172)

y-y1 = m(x-x1)
In this case y1 = 2
x1 = 1
m = slope = 3
y-2 = 3(x-1)
y-2 = 3x - 3
y = 3x -1 This equation is the answer
It shows us that the line has a slope of 3. If we assign zero to x we get y=-1 which means the line intersects the y axis at the following point
(0, -1)